Finn and Jake saw a carousel-a horizontal rotating platform- with radiuse when they decided to ride with it. As they accelerate, the movement of the carousel is at constant speed unit it has reached an angular velocity after 2 complete revolutions. What is the angular acceleration of the carousel during this time?

1
Expert's answer
2022-02-09T13:42:37-0500

we know that

Angular velocity


"w=2\pi n=2\pi\times\frac{2}{1}rad/sec=4\pi rad/sec"

Assume t=1 sec

Radius(r)=R

Linear velocity

"V=Rw"

Angular acceleration

"a_c=\frac{v^2}{R}"


"a_c={w^2}{R}=({4\pi})^2R=16\pi^2Rrad/sec^2"